Solve the quadratic equation. 0.2x^2-0.7x+0.2=0
Answer #1

Please see attached solution

Please like it0.24²_0.74 +0.2=0 2 x?- 7 x + = 0 =0 Qu² - 70+ 2 --btJ6ual (quadratic formo a=2 b=-70=2 -(-7) ŁJE7-4x2x2 x - 2x2 = ? 1 Jug 16